OpenStudy (anonymous):

What is the exact value of the square root of 24 over 2 the square root of 64.? Simplify if possible.

OpenStudy (anonymous):

\[\sqrt{24} \over 2\sqrt{64}\] is what the original equation looks like

OpenStudy (ankit042):

64=8*8 24 = 8*3 now you can cancel terms and solve

OpenStudy (mayankdevnani):

\[\huge \frac{\sqrt{24}}{2\sqrt{64}}\] where \(\large \sqrt{64}=\sqrt{8^2}=8\) we get, \[\huge \frac{\sqrt{24}}{2 \times 8}\] \[\huge \frac{\sqrt{24}}{16}\]

OpenStudy (mayankdevnani):

now simplify it

OpenStudy (mayankdevnani):

@glacher

OpenStudy (anonymous):

dont do it hes trying to trick you

OpenStudy (anonymous):

okay... so it would be \[2\sqrt{6} \over 16\]

OpenStudy (anonymous):

its a trap

OpenStudy (mayankdevnani):

you are going good. @glacher

OpenStudy (anonymous):

thank you @mayankdevnani very much! im usally good at math... but im not to good at Algebra 1

OpenStudy (mayankdevnani):

now, \[\huge \frac{2\sqrt{6}}{16}\] and 2 cancel out 16 by 8, we get \[\huge \frac{\sqrt{6}}{8}\]

OpenStudy (mayankdevnani):

now simplify it

OpenStudy (skullpatrol):

How can you simplify any further? @mayankdevnani ?

OpenStudy (mayankdevnani):

we can write \(\large \sqrt{6}\) as approximate 2.44 now, plug this value , we get \[\huge \frac{\sqrt{6}}{8}=\frac{2.44}{8}=0.305\] This is our approximate answer.

OpenStudy (mayankdevnani):

This is how we can simplify it. @skullpatrol

OpenStudy (mayankdevnani):

hope you understand. @glacher

OpenStudy (anonymous):

I do very much!! @mayankdevnani

OpenStudy (mayankdevnani):

good...